Circular No. 7977

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2002fl, 2002fm, 2002fn, 2002fo, 2002fp, 2002fq, 2002fr

Further to IAUC 7971, J. Raux, Laboratoire de Physique Nucléaire et de Hautes Energies, University of Paris VI, reports the discovery of seven apparent supernovae by the Supernova Cosmology Project. The supernovae were found with the Canada-France-Hawaii Telescope (CFHT) during a ‘rolling search’ spanning Mar. 16–May 20, with reference images having been obtained at CFHT during Mar.–Apr. 2001 (limiting mag 25.6, S/N = 5). Each field was observed approximately every third night (except for a few weather gaps), making it possible to construct lightcurves with at least seven nights’ data for each supernova. Spectral types and redshifts (given to 0.001 for host-galaxy redshifts, to 0.01 for supernovae redshifts) were obtained using the following telescopes (and instruments) listed: Keck (+ ESI), 2002fl; Very Large Telescope (VLT) UT1 Antu (+ FORS1), 2002fp; VLT UT2 Kuyen (+ FORS2), 2002fm and 2002fr; Gemini (+ GMOS), 2002fn. Objects without spectroscopy were identified as supernovae using their lightcurves (cf. Dahlen and Goobar 2002, PASP 114, 284); these lightcurves exhibited clear brightening over ∼ 2 weeks for each object, and each showed a ‘kink’ in the fading portion of the curve that is characteristic of type-Ia supernovae in the I band. Some two dozen active galactic nuclei were also observed in these same fields, but they were all within 0′′.01 of their galaxy cores, and they exhibited sudden brightening/fading in their lightcurves. SN 2002fl was 4′′.0 from its galaxy’s core, and SN 2002fo was 3′′.4 from the center of its apparent host galaxy. SN 2002fn was only 0′′.01 from its galaxy’s center. The other supernovae were in the range 0′′.3–0′′.6 from their host galaxies’ centers.

SN  2002 UT  α_{2000}  δ_{2000}  I  z Type
2002fl Mar. 16 14^{h}01^{m}52.^{s}33 + 4^{o}39′23′′.2 21.9 0.26 Ia
2002fm  20 14 00 29.75 + 4 46 50.1 22.5 0.448 ?
2002fn Apr. 12 14 00 07.95 + 4 54 45.4 24.1 0.677 I
2002fo 12 14 01 29.83 + 5 14 20.7 24.7 ? ?
2002fp  12 14 02 18.40 + 4 47 05.9 24.0 0.352 ?
2002fq 16 14 00 51.20 + 4 33 26.2 22.7 ? ?
2002fr  17 14 00 46.40 + 4 33 41.4 22.4 0.31 I?
2002 September 25  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green